# Quantum Emission from Coupled Spin Pairs in Hexagonal Boron Nitride

### v1

@@ III Discussion
Figure 4: The optical pumping loop of CB+ ONVB. (a) Bright excitation (red) from ground state (GS) to excited state (ES) is localized on ONVB(--). Through internal conversion, the ES state can go to doublet IS CB(0) + ONVB(0). The doublet IS can transfer to quartet IS through intersystem crossing with energy separation $J>0$. The $\pm$ 1/2 and 3/2 spin sublevels are separated by zero field splitting in quartet. The black dot indicates spin mixing between doublet and quartet. The black circle indicate the population magnitude at ground state. (b) The spin sublevel splitting under magnetic field in doublet and quartet. The $J$ is --3000, 0 and 3000 MHz, accordingly. (c) The population ratio between $|D_{s},+1/2\rangle$ and $|D_{s},-1/2\rangle$ at dim state upon magnetic field.
@@
Fig. 4b shows the splitting of spin sublevels under magnetic field for various $J$, which can exhibit different magnitude and sign depending on the relative orientation of coupled spins as appears in our DFT calculation. Due to the zero-field splitting of triplet sublevels of ONVB  the $|Q_{s},\pm 1/2\rangle$ always lie low and the main spin mixing is from $|Q_{s},\pm 3/2\rangle$ and $|D_{s},\pm 1/2\rangle$. The mixing is largely due to $|D_{s},-1/2\rangle$ and $|Q_{s},+3/2\rangle$ when $J<0$, while the mixing is between $|D_{s},+1/2\rangle$ and $|Q_{s},-3/2\rangle$ when $J>0$. The different coefficient of spin sublevels in eigenstates indicates the distinguishable population and this population difference, or ODMR contrast, could be tuned through magnetic field. More specifically, the variation in mixing with the quartet states can result in different non-radiative relaxation rates, leading to hyperpolarization and ODMR contrast. For each finite magnitude of J-coupling, we observe a strong dependence of the ratio of mixing coefficients on the magnetic field, as illustrated in Fig. 4b. This difference is particularly pronounced when the applied field is set close to the avoided crossing, reaching a maximum value of approximately 1.4. Ultimately, this should translate into differences in the internal conversion rates to the ground state. It is important to note that for $J$=0, i.e., when two non-interacting defects are present, both doublet states mix equally with the quartet manifold at each magnetic field, resulting in no discernible difference in non-radiative decay.
@@
To further substantiate the role of spin mixing in mediating the ODMR signal, we calculated the internal conversion rates from the metastable state. We approach this process by using the excited state and MS quartet for $r1$, MS quartet and ground state for $r2$, separately. With short distance below 1 nm, the small energy difference between bright and dim state and strong electron-phonon coupling lead to fast internal conversion rate around $10^{10}$ MHz. Therefore the system relax back to ground state rapidly without detectable optical signal. The orbital overlap integral and electron-phonon coupling decay with increasing distance where the $R$ distant dependent value of the integrals may be approximated by the overlap of two Slater orbitals which goes with $e^{-R}$ Silver and Ruedenberg (1968). By the two configurations CB5 and CB7, we can extrapolate the electron-phonon coupling and conversion rate to longer distance that cannot be simulated with DFT, as discussed in Supplementary Note 2. The estimated conversion rate $r1$ can be 6-8 MHz with 18 Å separation.
@@
With around 2 nm separation, the present coupled spin pair model could explain the previously reported signature of the ODMR signal from ground state $S=1/2$ defects reported in hBN sheets and nanotubes Scholten et al. (2023); Gao et al. (2023c); Guo et al. (2023). We note that if the metastable state is sufficiently long-living to carry out spin operations (e.g., ST1 defect in diamond Lee et al. (2013)) then this spin pair model can be extended to the ODMR centers where the spin-polarization is observed in the metastable state. In this case, the ground state could be singlet from an $S=1/2$ acceptor and $S=1/2$ donor whereas the metastable states comprise single and triplet spin pairs. We note a recent work (Ref. Scholten et al. (2024)) during the preparation of our manuscript which observes this phenomenon without adding microscopic models for the underlying processes.
@@
In summary, we propose the DAP emission is responsible for most of the visible ZPL emission at around 2.0 eV. The variance in the ZPL wavelengths and optical lifetimes reported in the experiments can be explained by the type and the distance of the donor defect with respect to a key acceptor defect, ONVB. Our simulations reveal the ODMR mechanism associated with the spin pairs in certain DAP structures activated in the dim metastable states which is manifested as ODMR detection of the $S=1/2$ ground state at non-zero constant magnetic fields. DAPs in wide band gap semiconductors and materials are quite common structures and are responsible for optical emissions and ODMR signal for Kramers doublet spin systems. The general mechanisms proposed here is insightful to not only the defects engineering in hBN for quantum information processing but also the optoelectronic applications of defects in other semiconductors.
@@ IV Methods
In this paper, we carry out the density functional theory (DFT) calculations by the Vienna ab initio simulation package (VASP) code Kresse and Furthmüller (1996a, b) with plane wave basis set. The valence electrons and core part are described with projector augmented wave (PAW) potentials Blöchl (1994); Kresse and Joubert (1999). A $8\times 8$ two-layered supercell model was used to avoid the interactions between periodic images and this is sufficient to use the $\Gamma$-point sampling scheme. The interlayer vdW interaction was described with DFT-D3 method of Grimme  Grimme et al. (2010). With mixing parameter $\alpha=0.32$, the hybrid density functional of Heyd, Scuseria, and Ernzerhof (HSE)  Heyd et al. (2003) could yield experimental optical gap around 6 eV without renormalization energy included and this functional is used to optimize the geometry and calculate electronic structures. The convergence threshold for the forces was set to 0.01 eV/Å with cutoff energy at 450 eV. $\Delta$SCF method Gali et al. (2009) was used to calculate excited states. The band alignment is based on the core level energy of one atom far away from the defect center in defective and it in perfect supercell.

### v2

@@ III Discussion
Figure 4: The optical pumping loop of CB– ONVB. a Bright excitation (red) from ground state (GS) to excited state (ES) is localized on ONVB(--). Through internal conversion, the ES state can go to doublet metastable state of CB0 – ONVB0, i.e., $D_{s}$. The $D_{s}$ are separated by $J>0$ energy from the quartet metastable state of CB0 – ONVB0, i.e., $Q_{s}$. The $\pm$ 1/2 and 3/2 spin sublevels are separated by the zero-field-splitting in $Q_{s}$. The black dot indicates spin mixing between $D_{s}$ and $Q_{s}$. The black circle indicates the population magnitude in the ground state. b The spin sublevels under magnetic field in $D_{s}$ and $Q_{s}$ with $J$ values at $-3000$, 00 and $+3000$ MHz, accordingly. c The population ratio between $|D_{s},+1/2\rangle$ and $|D_{s},-1/2\rangle$ in the metastable state at the given magnetic fields.
@@
Fig. 4b shows the splitting of spin sublevels under magnetic field for various $J$ values, which can exhibit different magnitude and sign depending on the relative orientation of coupled spins as appears in our DFT calculation. Due to the zero-field splitting of the triplet sublevels of ONVB  the $|Q_{s},\pm 1/2\rangle$ always lie low and the main spin mixing is from $|Q_{s},\pm 3/2\rangle$ and $|D_{s},\pm 1/2\rangle$. The mixing is largely due to $|D_{s},-1/2\rangle$ and $|Q_{s},+3/2\rangle$ when $J<0$, while the mixing is between $|D_{s},+1/2\rangle$ and $|Q_{s},-3/2\rangle$ when $J>0$. The different coefficient in the eigenstates of the spin sublevels implies population difference of the spin states which leads to ODMR contrast that could be tuned through constant magnetic fields. More specifically, the variation in mixing with the quartet states can result in different non-radiative relaxation rates, leading to hyperpolarization and ODMR contrast. For each finite magnitude of J-coupling, we observe a strong dependence of the ratio of mixing coefficients on the magnetic field, as illustrated in Fig. 4b. This difference is particularly pronounced when the applied field is set close to the avoided crossing, reaching a maximum value of approximately 1.4. Ultimately, this should translate into differences in the internal conversion rates to the ground state. It is important to note that for $J=0$, i.e., when two non-interacting defects are present, both doublet states mix equally with the quartet manifold at each magnetic field, resulting in no discernible difference in non-radiative decay.
@@
To further substantiate the role of spin mixing in mediating the ODMR signal, we calculated the internal conversion rates towards and from the MS metastable state. We approach this process by using the excited state and MS $D_{s}$ for $r1$, MS $D_{s}$ and doublet ground state for $r2$, separately. With short distance below 1 nm, the small energy difference between the optically active doublet excited state and the dim metastable state and the strong electron-phonon coupling lead to fast internal conversion rate around $10^{10}$ MHz. Therefore the system relaxes back to ground state rapidly without detectable optical signal. The orbital overlap integral and the resulting electron-phonon coupling decay with increasing distance where the $R$ distant dependent value of the integrals may be approximated by the overlap of two Slater orbitals which goes with $e^{-R}$ Silver and Ruedenberg (1968). By the two configurations CB5 and CB7, we can extrapolate the electron-phonon coupling and conversion rate to longer $R$ distance between the donor and acceptor that cannot be directly computed with DFT, as discussed in Supplementary Note 2. The estimated conversion rate $r1$ is at 6-8 MHz with 18 Å separation which results in competitive radiative and non-radiative pathways from the optically active excited state of the CB– ONVB DAP defects. This means that the ZPL wavelength, quantum yield and brightness of these DAP defects depend on the actual distance of the acceptor and donor. It is intriguing that both weak radiative and non-radiative ($r3$ internal conversion) decay pathways link the metastable state to the doublet ground state (Supplementary Note 2) that closes the optical spin-polarization loop of these DAP defects.
@@
With around 2 nm separation between CB and ONVB, the present coupled spin pair model could explain the previously reported signature of ODMR signals from ground state $S=1/2$ defects in hBN sheets and nanotubes Scholten et al. (2024); Gao et al. (2023); Guo et al. (2023). We underline that if the metastable state is sufficiently long-living to carry out spin operations (e.g., ST1 defect in diamond Lee et al. (2013)) then this spin-pair model can be extended to the ODMR centers in hBN where the spin-polarization is observed in the metastable state. In this case, the ground state could be singlet from an $S=1/2$ acceptor and $S=1/2$ donor whereas the metastable states comprise single and triplet spin pairs. We note recent works (Refs. Patel et al. (2024); Scholten et al. (2024); Robertson et al. (2024)) during the preparation of our manuscript which observe this phenomenon without adding microscopic models for the underlying processes.
@@
In summary, we propose that the DAP emission is responsible for most of the visible ZPL emission at around 2.0 eV. The variance in the ZPL wavelengths and optical lifetimes reported in the experiments can be explained by the type and the distance of the donor defect with respect to a key acceptor defect, ONVB. Our simulations reveal the ODMR mechanism associated with the spin pairs in certain DAP structures activated within the dim metastable states which is manifested as ODMR detection of the $S=1/2$ ground state at non-zero constant magnetic fields. Our model explains the challenge to produce indistinguishable SPEs in hBN as the location of the defects pairs should be well engineered. DAPs in wide band gap semiconductors and materials are quite common structures and are responsible for optical emissions and ODMR signal for Kramers doublet spin systems. The general mechanisms proposed here is insightful to not only the defects engineering in hBN for quantum information processing but also the optoelectronic applications of defects in other semiconductors.
@@ IV Methods
In this paper, we carry out the density functional theory (DFT) calculations by the Vienna ab initio simulation package (VASP) code Kresse and Furthmüller (1996a, b) with plane wave basis set. We applied a plane wave cutoff energy at 450 eV. The valence electrons and the core part are described with projector augmented wave (PAW) potentials Blöchl (1994); Kresse and Joubert (1999). A $8\times 8$ two-layered supercell model was used to avoid the interactions between periodic images and this is sufficient to use the $\Gamma$-point sampling scheme. The interlayer vdW interaction was described with DFT-D3 method of Grimme Grimme et al. (2010). With mixing parameter $\alpha=0.32$, the hybrid density functional of Heyd, Scuseria, and Ernzerhof (HSE) Heyd et al. (2003) could yield experimental optical gap around 6 eV without electron-phonon renormalization energy included and this functional is used to optimize the geometry and calculate electronic structures. The convergence threshold for the forces was set to 0.01 eV/Å. $\Delta$SCF method Gali et al. (2009) was used to calculate electronic excited states. The band alignment in the charge correction procedure is based on the core level energy of one selected atom far away from the defect center in the defective and the perfect supercell. The zero-field-splitting due to dipolar electron-spin electron-spin interaction was calculated within the PAW formalism Bodrog and Gali (2014) as implemented in VASP by Martijn Marsman. We applied the spin decontemination method to arrive at the final result Biktagirov et al. (2020).